Find the equation of each of these straight lines in the form of y=mx+c.
b) perpendicular to the line 3x+4y=5 passing through (3,-2)

Answer by mananth(16946)

3 x + 4 y = 5
Find the slope of this line

4 y = -3 x + 5
Divide by 4
y = - 3/ 4 x + 1 1/4
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 3/4

The slope of a line perpendicular to the above line will be the negative reciprocal 1 1/3
Because m1*m2 =-1
The slope of the required line will be 1 1/3

m= 1 1/3 ,point ( 3 , -2 )
Find b by plugging the values of m & the point in
y=mx+b
-2 = 4 + b
b= -6
m= 1 1/3
The required equation is y =4/ 3 x -6
m.ananth@hotmail.ca